# ParetoFlow: Guided Flows in Multi-Objective Optimization

### v1

@@ 1 Introduction
Offline optimization, a fundamental challenge in science and engineering, involves minimizing a black-box function using solely an offline dataset, with diverse applications ranging from protein design Sarkisyan et al. (2016); Angermueller et al. (2019) to neural architecture design Lu et al. (2023). Previous research primarily focuses on single-objective optimization, aiming to optimize a single desired property Trabucco et al. (2022); however, this fails to capture the complexities of real-world challenges that often require balancing multiple conflicting objectives, such as designing a neural architecture that demands both high accuracy and minimal parameter count Lu et al. (2023). In this study, we explore offline multi-objective optimization (MOO), leveraging an offline dataset of designs and their associated labels to simultaneously minimize multiple objectives.
@@ 3.1 Multi-Objective predictor guidance
Predictor Guidance. Originally, classifier guidance was proposed to direct sample generation toward specific image categories Dhariwal & Nichol (2021). This concept has been adapted for regression settings to guide molecule generation Lee et al. (2023); Jian et al. (2024). In this paper, we term this technique predictor guidance for a generalization. Based on Lemma 1 in Zheng et al. (2023), we derive predictor guidance in flow matching as:
@@ 4.4 Results and Analysis
$\mathcal{D}$(best) $16.82$$\pm$ $6.28$ $14.42$$\pm$ $4.11$ $15.00$$\pm$ $4.00$ $13.75$$\pm$ $6.91$ $18.06$$\pm$ $3.93$ $16.02$$\pm$ $5.13$
E2222E $10.91$$\pm$ $8.20$ $6.05$$\pm$ $3.32$ $12.50$$\pm$ $1.50$ $9.75$$\pm$ $4.97$ $9.69$$\pm$ $5.65$ $8.73$$\pm$ $5.88$
E2222E + GradNorm $12.73$$\pm$ $6.57$ $13.42$$\pm$ $5.54$ $8.50$$\pm$ $0.50$ $13.50$$\pm$ $5.12$ $14.19$$\pm$ $5.87$ $13.33$$\pm$ $5.84$
E2222E + PcGrad $9.55$$\pm$ $6.27$ $6.42$$\pm$ $3.18$ $16.50$$\pm$ $2.50$ $14.00$$\pm$ $3.16$ $10.88$$\pm$ $6.17$ $9.42$$\pm$ $5.68$
@@
MM $\underline{5.00}\pm\underline{4.18}$ $6.74$$\pm$ $3.81$ $16.50$$\pm$ $1.50$ $6.75$$\pm$ $4.32$ $6.69$$\pm$ $3.46$ $\underline{6.73}$$\pm$ $\underline{4.30}$
@@
MM + RoMA $13.36$$\pm$ $7.39$ $8.21$$\pm$ $5.75$ $10.00$$\pm$ $3.00$ $12.00$$\pm$ $2.45$ $10.25$$\pm$ $5.14$ $10.29$$\pm$ $6.03$
MM + IOM $7.00$$\pm$ $3.77$ $5.37$$\pm$ $3.60$ $6.50$$\pm$ $0.50$ $10.75$$\pm$ $1.92$ $7.25$$\pm$ $4.02$ $6.75$$\pm$ $3.88$
MM + ICT $14.55$$\pm$ $5.71$ $8.53$$\pm$ $3.12$ $9.50$$\pm$ $3.50$ $12.50$$\pm$ $7.12$ $11.75$$\pm$ $6.54$ $11.13$$\pm$ $5.77$
MM + Tri-Mentor $11.09$$\pm$ $5.84$ $9.05$$\pm$ $5.71$ $10.50$$\pm$ $1.50$ $13.00$$\pm$ $3.54$ $10.50$$\pm$ $5.82$ $10.29$$\pm$ $5.64$
MOEA/D + MM $10.64$$\pm$ $4.79$ $12.58$$\pm$ $5.02$ $11.00$$\pm$ $1.00$ $10.75$$\pm$ $6.87$ $12.12$$\pm$ $6.62$ $11.83$$\pm$ $5.65$
MOBO $11.00$$\pm$ $4.37$ $14.74$$\pm$ $3.82$ $17.00$$\pm$ $0.00$ $8.25$$\pm$ $6.61$ $11.00$$\pm$ $5.79$ $12.38$$\pm$ $5.31$
MOBO-$q$ParEGO $13.45$$\pm$ $3.94$ $16.63$$\pm$ $3.77$ $21.00$$\pm$ $0.00$ $12.75$$\pm$ $8.04$ $17.69$$\pm$ $4.55$ $16.15$$\pm$ $4.90$
@@ A.4 Training Details
We adopt the predictor training configurations from Xue et al. (2024), utilizing a multiple model setup. Each predictor consists of a $3$-layer MLP with ReLU activations, featuring a hidden layer size of $2048$. These models are trained over $200$ epochs with a batch size of $128$, using the Adam optimizer Kingma (2014) at an initial learning rate of $1\times 10^{-3}$, and a decay rate of $0.98$ per epoch. Flow matching training follows protocols from Tomczak (2022), employing a $4$-layer MLP with SeLU activations and a hidden layer size of $512$. The model undergoes $1000$ training epochs with early stopping after $20$ epochs of no improvement, with a batch size of $128$ and the Adam optimizer.
@@ A.9 Effectiveness of local filtering
To further verify the effectiveness of local filtering, we conduct experiments on the convex-PF task ZDT1111 and the nonconvex-PF task ZDT2222. When we remove the local filtering, the performance of ZDT1111 nearly does not change: from $4.88\pm 0.02$ to $4.81\pm 0.08$. In contrast, the performance on ZDT2222 drops obviously: from $6.79\pm 0.16$ to $5.78\pm 0.15$. This demonstrates the effectiveness of our local filtering scheme in handling nonconvex PFs and verifies its underlying motivation.

### v2

@@ 1 Introduction
Offline optimization, a fundamental challenge in science and engineering, involves minimizing a black-box function using solely an offline dataset, with diverse applications ranging from protein design Sarkisyan et al. (2016); Angermüller et al. (2020) to neural architecture design Lu et al. (2023). Previous research primarily focuses on single-objective optimization, aiming to optimize a single desired property Trabucco et al. (2022); however, this fails to capture the complexities of real-world challenges that often require balancing multiple conflicting objectives, such as designing a neural architecture that demands both high accuracy and minimal parameter count Lu et al. (2023). In this study, we explore offline multi-objective optimization (MOO), leveraging an offline dataset of designs and their associated labels to simultaneously minimize multiple objectives.
@@ 3.1 Multi-Objective predictor guidance
Predictor Guidance. Originally, classifier guidance was proposed to direct sample generation toward specific image categories Dhariwal & Nichol (2021). This concept has been adapted for regression settings to guide molecule generation Lee et al. (2023); Jian et al. (2024); Chen et al. (2025). In this paper, we term this technique predictor guidance for a generalization. Based on Lemma 1 in Zheng et al. (2023), we derive predictor guidance in flow matching as:
@@ 4.4 Results and Analysis
D-Best $16.82$$\pm$ $6.28$ $14.42$$\pm$ $4.11$ $15.00$$\pm$ $4.00$ $13.75$$\pm$ $6.91$ $18.06$$\pm$ $3.93$ $16.02$$\pm$ $5.13$
E2E $10.91$$\pm$ $8.20$ $6.05$$\pm$ $3.32$ $12.50$$\pm$ $1.50$ $9.75$$\pm$ $4.97$ $9.69$$\pm$ $5.65$ $8.73$$\pm$ $5.88$
E2E + GradNorm $12.64$$\pm$ $6.68$ $13.42$$\pm$ $5.54$ $8.50$$\pm$ $0.50$ $13.50$$\pm$ $5.12$ $14.19$$\pm$ $5.87$ $13.31$$\pm$ $5.87$
E2E + PcGrad $9.45$$\pm$ $6.37$ $6.42$$\pm$ $3.18$ $16.50$$\pm$ $2.50$ $14.00$$\pm$ $3.16$ $10.88$$\pm$ $6.17$ $9.40$$\pm$ $5.70$
@@
MM $\underline{4.91}$$\pm$ $\underline{4.17}$ $6.74$$\pm$ $3.81$ $16.50$$\pm$ $1.50$ $6.75$$\pm$ $4.32$ $6.69$$\pm$ $3.46$ $\underline{6.71}$$\pm$ $\underline{4.31}$
@@
MM + RoMA $13.27$$\pm$ $7.53$ $8.21$$\pm$ $5.75$ $10.00$$\pm$ $3.00$ $12.00$$\pm$ $2.45$ $10.25$$\pm$ $5.14$ $10.27$$\pm$ $6.06$
MM + IOM $6.91$$\pm$ $3.78$ $5.37$$\pm$ $3.60$ $6.50$$\pm$ $0.50$ $10.75$$\pm$ $1.92$ $7.25$$\pm$ $4.02$ $6.73$$\pm$ $3.88$
MM + ICT $14.45$$\pm$ $5.77$ $8.53$$\pm$ $3.12$ $9.50$$\pm$ $3.50$ $12.50$$\pm$ $7.12$ $11.75$$\pm$ $6.54$ $11.12$$\pm$ $5.77$
MM + Tri-Mentor $11.00$$\pm$ $5.89$ $9.05$$\pm$ $5.71$ $10.50$$\pm$ $1.50$ $13.00$$\pm$ $3.54$ $10.50$$\pm$ $5.82$ $10.27$$\pm$ $5.65$
MOEA/D + MM $10.55$$\pm$ $4.83$ $12.58$$\pm$ $5.02$ $11.00$$\pm$ $1.00$ $10.75$$\pm$ $6.87$ $12.12$$\pm$ $6.62$ $11.81$$\pm$ $5.66$
MOBO $10.91$$\pm$ $4.42$ $14.74$$\pm$ $3.82$ $17.00$$\pm$ $0.00$ $8.25$$\pm$ $6.61$ $11.00$$\pm$ $5.79$ $12.37$$\pm$ $5.32$
MOBO-$q$ParEGO $13.36$$\pm$ $3.98$ $16.63$$\pm$ $3.77$ $21.00$$\pm$ $0.00$ $12.75$$\pm$ $8.04$ $17.69$$\pm$ $4.55$ $16.13$$\pm$ $4.91$
@@ 6 Conclusion
7 Acknowledgements
This research was partially funded by the Fonds de recherche du Québec – Nature et technologies. We also gratefully acknowledge CIFAR for its support through the AI Chairs program.
We thank Mattie Tesfaldet and Alexander Tong from Mila, along with Chin-Wei Huang from Microsoft Research, for their insightful discussions on score-based models. We further appreciate Jiarui Lu from Mila for his valuable suggestions regarding the presentation of this paper.
@@ A.4 Training Details
We adopt the predictor training configurations from Xue et al. (2024), utilizing a multiple model setup. Each predictor consists of a $3$-layer MLP with ReLU activations, featuring a hidden layer size of $2048$. These models are trained over $200$ epochs with a batch size of $128$, using the Adam optimizer Kingma & Ba (2015) at an initial learning rate of $1\times 10^{-3}$, and a decay rate of $0.98$ per epoch. Flow matching training follows protocols from Tomczak (2022), employing a $4$-layer MLP with SeLU activations and a hidden layer size of $512$. The model undergoes $1000$ training epochs with early stopping after $20$ epochs of no improvement, with a batch size of $128$ and the Adam optimizer.
@@ A.9 Effectiveness of local filtering
To further verify the effectiveness of local filtering, we conduct experiments on the convex-PF task ZDT1111 and the nonconvex-PF task ZDT2222. When we remove the local filtering, the performance of ZDT1111 nearly does not change: from $4.30\pm 0.02$ to $4.29\pm 0.04$. In contrast, the performance on ZDT2222 drops obviously: from $6.79\pm 0.16$ to $5.78\pm 0.15$. This demonstrates the effectiveness of our local filtering scheme in handling nonconvex PFs and verifies its underlying motivation.